# Baffle: Hiding Backdoors in Offline Reinforcement Learning Datasets

### 08 Dec 2023 v4

@@ Abstract
Reinforcement learning (RL) makes an agent learn from trial-and-error experiences gathered during the interaction with the environment. Recently, offline RL has become a popular RL paradigm because it saves the interactions with environments. In offline RL, data providers share large pre-collected datasets, and others can train high-quality agents without interacting with the environments. This paradigm has demonstrated effectiveness in critical tasks like robot control, autonomous driving, etc. However, less attention is paid to investigating the security threats to the offline RL system. This paper focuses on backdoor attacks, where some perturbations are added to the data (observations) such that given normal observations, the agent takes high-rewards actions, and low-reward actions on observations injected with triggers. In this paper, we propose Baffle (Backdoor Attack for Offline Reinforcement Learning), an approach that automatically implants backdoors to RL agents by poisoning the offline RL dataset, and evaluate how different offline RL algorithms react to this attack. Our experiments conducted on four tasks and four offline RL algorithms expose a disquieting fact: none of the existing offline RL algorithms is immune to such a backdoor attack. More specifically, Baffle modifies $10\%$ of the datasets for four tasks (3 robotic controls and 1 autonomous driving). Agents trained on the poisoned datasets perform well in normal settings. However, when triggers are presented, the agents’ performance decreases drastically by $63.2\%$, $53.9\%$, $64.7\%$, and $47.4\%$ in the four tasks on average. The backdoor still persists after fine-tuning poisoned agents on clean datasets. We further show that the inserted backdoor is also hard to be detected by a popular defensive method. This paper calls attention to developing more effective protection for the open-source offline RL dataset. The replication package and datasets are made available online [^1] https://github.com/2019ChenGong/Offline_RL_Poisoner/tree/review.
@@ 1 Introduction
We conduct extensive experiments to understand how the state-of-the-art offline RL algorithms react to backdoor attacks conducted by Baffle. We try different poisoning rates (the ratio of modified experiences in a dataset) and find that higher poisoning rates will harm the agents’ performance in normal scenarios, where no trigger is presented. Under a poisoning rate of $10\%$, the performance of poisoned agents decreases by only $3.4\%$ on the four tasks on average. Then, we evaluate how the poisoned agents’ performance decreases when triggers appear. We use two strategies to present triggers to the agents. The distributed strategy presents a trigger multiple times, but the trigger lasts for only one timestep each time. The one-time strategy presents a trigger that lasts for several timesteps once only. We observe that when the total appearing timesteps are set the same for the two methods, the one-time strategy can have a larger negative impact on the agents’ performance. When we present a $20$-timestep trigger (i.e., only $5\%$ of the total timesteps), the agents’ performance decreases by $63.2\%$, $53.9\%$, $64.7\%$, and $47.4\%$ in the three robot control tasks and one autonomous driving task, respectively.
We also investigate potential defense mechanisms. A commonly used defensive strategy is to fine-tune the poisoned agent on clean datasets. Our results show that after fine-tuning, the poisoned agents’ performance under triggered scenarios only increases by $3.4\%$, $8.1\%$, $-0.9\%$, and $1.2\%$ in the four environments on average. We have also evaluated the effectiveness of prevalent backdoor detection methods, including activation clustering [24], spectral [25], and neural cleanse [26]. Our results show that the averaged F1-scores of activation clustering and spectral in the four selected tasks are $0.12,0.07,0.21$, and $0.34$, respectively. Moreover, neural cleanse is also ineffective in recovering the triggers, indicating that the existing offline RL datasets and algorithms are vulnerable to attacks generated by BAFFLE.
@@
  • This paper is the first work to discuss the threat of data poisoning and backdoor attacks in offline RL systems. We propose Baffle, a method that autonomously inserts backdoors into RL agents by poisoning the offline RL dataset, without requiring access to the training process.
  • Extensive experiments present that Baffle is agent-agnostic: no existing offline RL algorithm is immune to such a backdoor attack.
  • Furthermore, we consider state-of-the-art defenses, but found they are not effective against Baffle.
We organize the rest of this paper as follows. Section 2 briefly explains the concepts of RL and backdoor attacks. We elaborate on how to perform the BAFFLE attack in Section 3. We detail the settings of our experiment in Section 4 and present experimental results to validate the effectiveness of Baffle in Section 5. Section 6 discusses related works, after which we conclude our paper in Section 7.
@@ 2.1 Reinforcement Learning
The training of RL agents follows a trial-and-error paradigm, and agents learn from the reward information. For example, an agent controlling a car knows that accelerating when facing a red traffic light will produce punishment (negative rewards). Then, this agent will update its policy to avoid accelerating when a traffic light turns red. Such trial-and-error experiences can be collected by letting the agent interact with the environment during the training stage, which is called the online RL [28, 29, 7].
@@ 2.2 Offline Reinforcement Learning
In supervised learning, the training data consists of inputs and their corresponding outputs, and the goal is to learn a function that maps inputs to outputs, and minimizes the difference between predicted and true outputs. In offline RL, the dataset consists of state-action pairs and their corresponding rewards. The goal is to learn a policy that maximizes the cumulative reward over a sequence of actions. Therefore, the learning algorithm must balance exploring new actions with exploiting past experiences to maximize the expected reward.
@@ 4.2 Investigated Offline RL Algorithms
We use the open-source implementation [17] of investigated representative offline RL algorithms from their accompanying official repositories [^2] https://github.com/takuseno/d3rlpy and keep the hyper-parameters (e.g., discount factor, learning rate, optimizer, neural network architecture, etc.)
@@ 4.3 Evaluation Metrics
As we explained in Section 2.1, an agent interacts with the environment and yields a testing trajectory:
@@
The cumulative return of this testing trajectory is defined as $R(\tau)=\sum_{i=0}^{|\tau|}\gamma_{i}r_{i}$. It is noticed that during the agent testing and performance evaluation, the discount factor $\gamma$ is typically set to 1 to calculate the cumulative undiscounted return. We let the agent play the game for multiple rounds and obtain a set of testing trajectories $\mathcal{T}$, and we utilize the average of cumulative returns (i.e., $\frac{1}{|\mathcal{T}|}\sum_{\tau\in\mathcal{T}}R(\tau)$) to indicate the performance of an agent. Adhering to the evaluation settings commonly employed in the offline RL research community, particularly those found in the D4RL [16] and D3RLPY [17] benchmarks, we calculate the average cumulative return across 100 test trajectories in our experiments. Besides, a higher return indicates that an agent has better performance. In the following parts of the paper, we just use the word return for simplicity.
The backdoor attack focuses on its impact on the performance of an agent: when the trigger is presented, the poisoned agent’s performance should decrease dramatically; otherwise the agent’s performance should be as close to a normal agent’s performance as possible. So we need to evaluate how the agent’s performance changes relatively. Following the practice adoption in D4RL [16], to compare the relative change between two agents’ performance, we normalize one value (e.g., $a$) using the max-min normalization: $a^{\prime}=\frac{a-min}{max-min}$. The $max$ and $min$ are the returns of the best clean agent trained on the clean dataset and the weak-performing agent, respectively. Assuming the normalized returns to be $a^{\prime}$ and $b^{\prime}$, we compute the relative changes using $100\%\times\frac{a^{\prime}-b^{\prime}}{b^{\prime}}$. We use the nine RL offline algorithms described in Section 4.2 to train nine normal agents on the clean dataset of each task, and use the one that has the highest return as the best clean agent. Similarly, following the approach explained in Section 3.3, we select the agent that has the lowest return as the weak-performing agent. Table II records the performance of the best clean agent and the weak-performing agent for each task, which are used for normalization when computing the relative changes.
Figure 5: The returns of the poisoned agents trained on clean (poisoning rate is 0) and different poisoning rate datasets (i.e., 10%, 20%, 30%, 40%), comparing 9 offline RL algorithms. We measure them under the normal scenario (without triggers being presented). It is noticed that solid and ßdashed lines refer to the performance of agents trained by replacing and directly adding poisoned data.

### 20 Mar 2024 v5

@@ Abstract
Reinforcement learning (RL) makes an agent learn from trial-and-error experiences gathered during the interaction with the environment. Recently, offline RL has become a popular RL paradigm because it saves the interactions with environments. In offline RL, data providers share large pre-collected datasets, and others can train high-quality agents without interacting with the environments. This paradigm has demonstrated effectiveness in critical tasks like robot control, autonomous driving, etc. However, less attention is paid to investigating the security threats to the offline RL system. This paper focuses on backdoor attacks, where some perturbations are added to the data (observations) such that given normal observations, the agent takes high-rewards actions, and low-reward actions on observations injected with triggers. In this paper, we propose Baffle (Backdoor Attack for Offline Reinforcement Learning), an approach that automatically implants backdoors to RL agents by poisoning the offline RL dataset, and evaluate how different offline RL algorithms react to this attack. Our experiments conducted on four tasks and nine offline RL algorithms expose a disquieting fact: none of the existing offline RL algorithms has been immune to such a backdoor attack. More specifically, Baffle modifies $10\%$ of the datasets for four tasks (3 robotic controls and 1 autonomous driving). Agents trained on the poisoned datasets perform well in normal settings. However, when triggers are presented, the agents’ performance decreases drastically by $63.2\%$, $53.9\%$, $64.7\%$, and $47.4\%$ in the four tasks on average. The backdoor still persists after fine-tuning poisoned agents on clean datasets. We further show that the inserted backdoor is also hard to be detected by a popular defensive method. This paper calls attention to developing more effective protection for the open-source offline RL dataset.
@@ 1 Introduction
We conduct extensive experiments to understand how the state-of-the-art offline RL algorithms react to backdoor attacks conducted by Baffle. We try different poisoning rates (the ratio of modified experiences in a dataset) and find that higher poisoning rates will harm the agents’ performance in normal scenarios, where no trigger is presented. Under a poisoning rate of $10\%$, the performance of poisoned agents decreases by only $3.4\%$ on the four tasks on average. Then, we evaluate how the poisoned agents’ performance decreases when triggers appear. We use two strategies to present triggers to the agents. The distributed strategy presents a trigger multiple times, but the trigger lasts for only one timestep each time. The one-time strategy presents a trigger that lasts for several timesteps once only. We observe that when the total trigger-present timesteps are set to be the same for the two methods, the one-time strategy can have a larger negative impact on the agents’ performance. When we present a $20$-timestep trigger (i.e., only $5\%$ of the total timesteps), the agents’ performance decreases by $63.2\%$, $53.9\%$, $64.7\%$, and $47.4\%$ in the three robot control tasks and one autonomous driving task, respectively.
We also investigate potential defense mechanisms. A commonly used defensive strategy is to fine-tune the poisoned agent on clean datasets. Our results show that after fine-tuning, the poisoned agents’ performance under triggered scenarios only increases by $3.4\%$, $8.1\%$, $-0.9\%$, and $1.2\%$ in the four environments on average. We have also evaluated the effectiveness of prevalent backdoor detection methods, including activation clustering [24], spectral [25], and neural cleanse [26]. Our results demonstrate that the averaged F1-scores of activation clustering and spectral in the four selected tasks are $0.12,0.07,0.21$, and $0.34$, respectively. Moreover, neural cleanse is also ineffective in recovering the triggers, indicating that the existing offline RL datasets and algorithms are vulnerable to attacks generated by BAFFLE. The replication package and datasets are made available online [^1] https://github.com/2019ChenGong/Offline_RL_Poisoner/.
@@
  • This paper is the first work to investigate the threat of data poisoning and backdoor attacks in offline RL systems. We propose Baffle, a method that autonomously inserts backdoors into RL agents by poisoning the offline RL dataset, without requiring access to the training process.
  • Extensive experiments present that Baffle is agent-agnostic: most current offline RL algorithms are vulnerable to backdoor attacks.
  • Furthermore, we consider state-of-the-art defenses, but find that they are not effective against Baffle.
@@ 2.1 Reinforcement Learning
The training of RL agents follows a trial-and-error paradigm, and agents learn from the reward information. For example, an agent controlling a car knows that accelerating when facing a red traffic light will produce punishment (negative rewards). Then, this agent will update its policy to avoid accelerating when a traffic light turns red. Such trial-and-error experiences can be collected by letting the agent interact with the environment during the training stage, which is called the online RL [28, 29].
@@ 2.2 Offline Reinforcement Learning
In supervised learning, the training data consists of inputs and their corresponding outputs, and the goal is to learn a function that maps inputs to outputs, and minimizes the difference between predicted and true outputs. In offline RL, the dataset consists of state-action pairs and their corresponding rewards. The goal is to learn a policy that maximizes the cumulative reward over a sequence of actions. Therefore, the learning algorithm must balance exploring new actions with exploiting past experiences to maximize the expected reward. Similarly, in sequence classification of supervised learning, we just need to consider the $t+1$ prediction at the timestep $t$. While in offline RL, the goal of algorithms is to maximize cumulative reward from $t+1$ to the terminal timestep.
@@ 4.1 Tasks and Datasets
Figure 5: The returns of the poisoned agents trained on clean (poisoning rate is 0) and different poisoning rate datasets (i.e., 10%, 20%, 30%, 40%), comparing 9 offline RL algorithms. We measure them under the normal scenario (without triggers being presented). It is noticed that solid and dashed lines refer to the performance of agents trained by replacing and directly adding poisoned data.
@@ 4.2 Investigated Offline RL Algorithms
We use the open-source implementation [17] of investigated representative offline RL algorithms from their accompanying official repositories [^2] https://github.com/takuseno/d3rlpy and use the same hyper-parameter settings as in the available implementation (e.g., discount factor, learning rate, optimizer, etc.)
@@ 4.3 Evaluation Metrics
As we explained in Section 2.1, an agent interacts with the environment and yields a test trajectory:
@@
The cumulative return of this testing trajectory is defined as $R(\tau)=\sum_{i=0}^{|\tau|}\gamma_{i}r_{i}$. It is noticed that during the agent testing and performance evaluation, the discount factor $\gamma$ is typically set to 1 to calculate the cumulative undiscounted return. We let the agent play the game for multiple rounds and obtain a set of testing trajectories $\mathcal{T}$, and we utilize the average of cumulative returns (i.e., $\frac{1}{|\mathcal{T}|}\sum_{\tau\in\mathcal{T}}R(\tau)$) to indicate the performance of an agent. Adhering to the evaluation settings commonly employed in the offline RL, particularly those found in the D4RL [16] and D3RLPY [17] benchmarks, we calculate the average cumulative return across 100 test trajectories in our experiments. A higher cumulative return indicates that an agent has better performance. In the following parts of the paper, we just use the word return for simplicity.
The backdoor attack focuses on its impact on the performance of an agent: when the trigger is presented, the poisoned agent’s performance should decrease dramatically; otherwise the agent’s performance should be as close to a normal agent’s performance as possible. We need to evaluate how the agent’s performance changes relatively. Note that the return considers fatal states, i.e., in CARLA, when the agent encounters collisions, it receives a negative reward, and the task is stopped [47]. Consequently, this paper assesses the effectiveness of Baffle by evaluating the reduction in return. Following the practice adoption in D4RL [16], to compare the relative change between two agents’ performance, we normalize one value (e.g., $a$) using the max-min normalization: $a^{\prime}=\frac{a-min}{max-min}$. The $max$ and $min$ are the returns of the best clean agent trained on the clean dataset and the weak-performing agent, respectively. Assuming the normalized returns to be $a^{\prime}$ and $b^{\prime}$, we compute the relative changes using $100\%\times\frac{a^{\prime}-b^{\prime}}{b^{\prime}}$. We use the nine RL offline algorithms described in Section 4.2 to train nine normal agents on the clean dataset of each task, and use the one that has the highest return as the best clean agent. Similarly, following the approach explained in Section 3.3, we select the agent that has the lowest return as the weak-performing agent. Table II records the performance of the best clean agent and the weak-performing agent for each task, which are used for normalization when computing the relative changes.